### v1

@@ 4.2 LLM Evaluation on CASTLE Benchmark
On the CASTLE dataset, LLMs exhibited notably strong performance. In particular, GPT-o3-Mini achieved the highest overall score of 977 points, correctly identifying 126 out of the 150 known vulnerabilities. When examining the true positives across different LLM variants, we observed that GPT-4o and GPT-4o-Mini generated a higher count of accurate detections than GPT-o1 or GPT-o3-Mini for true positives. However, the reasoning-oriented models consistently produced fewer false positives, suggesting that their internal steps for “validating” potential vulnerabilities lead to more precise outcomes.
Our findings indicate that modern LLMs are adept at pinpointing vulnerabilities in short, self-contained C programs. We conjecture that their neural architectures confer an inherent advantage in pattern recognition, while more advanced “reasoning” models appear better at avoiding out false detections. As a result, LLM-based approaches rival—and often surpass—several classical static analysis tools in detecting common software flaws within compact code segments. However, the next section highlights several limitations and issues for LLMs.
@@ 5 Limitations
Microbenchmark Scope: A fundamental concern with any microbenchmark-based study is its limited scope. Although the CASTLE dataset covers 25 distinct CWEs, each test typically focuses on a single vulnerability in an isolated context. Real-world software often exhibits multi-faceted security flaws spanning tens or hundreds of files. Consequently, tools optimized for detecting specific vulnerabilities may perform artificially well on microbenchmarks while missing complex, cross-file weaknesses that only arise in large-scale applications. Regardless, tools did not perform that well on even this small test, indicating that their high false positive rates would be a problem for longer contexts.
Lack of Large Code Samples: Preliminary testing with a synthetic 400+ line C program created by merging multiple non-vulnerable test cases, revealed that LLMs exhibit a higher tendency to report false positives when dealing with larger codebases. Similarly, when one hidden vulnerability was introduced into this combined file, most LLMs failed to detect it reliably, suggesting that these models’ effectiveness may taper off with increasing code length. Formal verification approaches also suffer from scalability issues, such as state explosion, and may require lowered bounds that reduce their thoroughness. By contrast, classical static application security testers (SAST) can handle extensive projects more efficiently, yet their propensity for false positives undercuts overall usefulness in large-scale deployments.
@@ 6 Conclusion
In this study, we introduced the CASTLE benchmark, a curated collection of 250 compilable C micro-benchmarks covering 25 major CWEs, and proposed the CASTLE Score to evaluate a diverse set of vulnerability detection tools, including static analyzers, formal verification methods, and large language models (LLMs). Our work aimed to address the following research questions:
@@
    Answer: LLMs exhibit high effectiveness on compact code snippets, with GPT-o3-Mini scoring the highest (977 points) by identifying 126 out of 150 vulnerabilities. However, their performance declines on larger codebases, where false positives increase and hidden vulnerabilities often remain undetected. Static analyzers demonstrate moderate performance but produce numerous false positives, creating substantial manual triage overhead. Formal verification tools yield minimal false positives within their supported classes (e.g., memory safety) but cannot detect certain higher-level vulnerabilities such as SQL injection, limiting their coverage.
@@
    Answer: Tool combinations frequently outperform individual tools, particularly when they offset each other’s weaknesses. For instance, ESBMC (formal verification) combined with CodeQL achieved the highest two-tool CASTLE Score (814). Although overlapping detections can inflate false positives, well-chosen pairs leverage complementary detection strategies, thereby enhancing overall reliability.
@@
Implications and Future Work. Although micro-benchmarks efficiently reveal how tools behave on targeted vulnerabilities, they may not reflect the full complexity of production-scale systems. Preliminary experiments show that LLMs and formal verification tools both face significant scalability barriers when analyzing large codebases. Ultimately, the insights gained through CASTLE underscore the importance of selecting and combining tools to fit specific project requirements, rather than relying on any single method for comprehensive security assurance.

### v2

@@ 4.2 LLM Evaluation on CASTLE Benchmark
On the CASTLE dataset, LLMs exhibited notably strong performance. In particular, GPT-o3-Mini achieved the highest overall score of 977 points, correctly identifying 126 out of the 150 known vulnerabilities. When examining the true positives across different LLM variants, we observed that GPT-4o and GPT-4o-Mini generated more accurate detections than GPT-o1 or GPT-o3-Mini for true positives. However, the reasoning-oriented models consistently produced fewer false positives, suggesting that their internal steps for “validating” potential vulnerabilities lead to more precise outcomes.
Our findings indicate that modern LLMs can pinpoint vulnerabilities in short, self-contained C programs. We conjecture that their neural architectures confer an inherent advantage in pattern recognition, while more advanced “reasoning” models appear better at avoiding out false detections. As a result, LLM-based approaches rival—and often surpass—several classical static analysis tools in detecting common software flaws within compact code segments. However, the next section highlights several limitations and issues for LLMs.
@@ 5 Limitations
Microbenchmark Scope: A fundamental concern with any microbenchmark-based study is its limited scope. Although the CASTLE dataset covers 25 distinct CWEs, each test typically focuses on a single vulnerability in an isolated context. Real-world software often exhibits multi-faceted security flaws spanning tens or hundreds of files. Consequently, tools optimized for detecting specific vulnerabilities may perform artificially well on microbenchmarks while missing complex, cross-file weaknesses that only arise in large-scale applications. Regardless, tools did not perform well on even this small test, indicating that their high false positive rates would be a problem for longer contexts.
Lack of Large Code Samples: Preliminary testing with a synthetic 400+ line C program created by merging multiple non-vulnerable test cases, revealed that LLMs tend to report false positives when dealing with larger codebases. Similarly, when one hidden vulnerability was introduced into this combined file, most LLMs failed to detect it reliably, suggesting that these models’ effectiveness may taper off with increasing code length. Formal verification approaches also suffer from scalability issues, such as state explosion, and may require lowered bounds that reduce their thoroughness. By contrast, classical static application security testers (SAST) can handle extensive projects more efficiently, yet their propensity for false positives undercuts overall usefulness in large-scale deployments.
@@ 6 Conclusion
In this study, we introduced the CASTLE benchmark, a curated collection of 250 compilable C micro-benchmarks covering 25 major CWEs. We proposed the CASTLE Score to evaluate diverse vulnerability detection tools, including static analyzers, formal verification methods, and large language models (LLMs). Our work aimed to address the following research questions:
@@
    Answer: LLMs exhibit high effectiveness on compact code snippets, with GPT-o3-Mini scoring the highest (977 points) by identifying 126 out of 150 vulnerabilities. However, their performance declines on larger codebases, where false positives increase and hidden vulnerabilities often remain undetected. Static analyzers perform moderately but produce numerous false positives, creating substantial manual triage overhead. Formal verification tools yield minimal false positives within their supported classes (e.g., memory safety) but cannot detect certain higher-level vulnerabilities such as SQL injection, limiting their coverage.
@@
    Answer: Tool combinations frequently outperform individual tools, particularly when they offset each other’s weaknesses. For instance, ESBMC (formal verification) combined with CodeQL achieved the highest two-tool CASTLE Score (814). Although overlapping detections can inflate false positives, well-chosen pairs leverage complementary detection strategies, enhancing overall reliability.
@@
Implications and Future Work. Although micro-benchmarks efficiently reveal how tools behave on targeted vulnerabilities, they may not reflect the full complexity of production-scale systems. Preliminary experiments show that LLMs and formal verification tools both face significant scalability barriers when analyzing large codebases. Ultimately, the insights gained through CASTLE underscore the importance of selecting and combining tools to fit specific project requirements rather than relying on any single method for comprehensive security assurance.